# Single-loop Projection-free and Projected Gradient-based Algorithms for Nonconvex-concave Saddle Point Problems with Bilevel Structure

### v1

@@ 4.3 Analysis of Gradients Estimation
In this section, we aim to assess how accurately the gradient estimates, denoted as $G^{x}_{k}$ and $G^{y}_{k}$, approximate the true gradients $\nabla_{x}\mathcal{L}(x_{k},y_{k})$ and $\nabla_{y}\mathcal{L}(x_{k},y_{k})$, respectively. Through Lemmas 4.8 and 4.9, we analyze the bounds on the error in estimating these gradients. We demonstrate that the errors may depend linearly on certain terms or involve consecutive terms, helping us in controlling the error. Later, in the convergence analysis, we will also specify the step-sizes to ensure convergence.
@@
where $\Gamma_{1},\Gamma_{3}$ are defined in Lemma 4.6 and 4.7, respectively, $\Gamma_{4}\triangleq D_{\mathcal{X}}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|(\frac{\tilde{C}_{2}}{1-\beta}+\frac{\rho\tilde{C}}{(1-\rho)^{2}})+\frac{\rho C_{\theta x}^{g}D_{\mathcal{X}}}{1-\rho}\left\|w_{0}-v(x_{0},y_{0})\right\|$ and $\Gamma_{5}\triangleq\frac{\gamma\beta\mathbf{L}_{\theta}}{1-\beta}(1+\frac{\tilde{C}}{1-\rho})+\frac{\rho\gamma\mathbf{C}_{v1}C_{\theta x}^{g}}{1-\rho}$. In particular, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$ and $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}\left(\gamma(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\gamma\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\gamma\kappa_{g}^{4}\right)$.
@@
$\quad-\left.\nabla_{\theta x}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})\right\|$ (48)
@@
$\leq L_{x\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+C_{\theta x}^{g}\left\|w_{k+1}-v(x_{k},y_{k})\right\|+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}\left\|\ \theta_{k}-\theta^{*}(x_{k})\right\|$ (49)
where the last inequality is obtained using Assumption 2.2 and the triangle inequality. Next, utilizing Lemma 4.1 and 4.2 we can further provide upper-bounds for the terms in RHS of the above inequality to obtain the desired result in (4.8).
@@
$\left\|\nabla_{y}\mathcal{L}(x_{k},y_{k})-G_{k}^{y}\right\|\leq\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\mu D_{\mathcal{Y}}.$ (50)
@@
$\leq L_{y\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+\mu D_{\mathcal{Y}},$ (51)
where in the last inequality we use Lipschitz continuity of $\nabla_{y}\Phi(x,\cdot,y)$ and boundedness of set $Y$. Next, utilizing Lemma 4.1 we can further provide upper-bounds for the term in RHS of the above inequality as follows
@@ 5.1 Convergence Rate of Algorithm 1
Suppose Assumptions 2.1, 2.2 hold and let $\{(x_{k},y_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 1 with step-sizes $\gamma_{k}=\gamma>0$, $\eta_{k}=\eta>0$, $\alpha>0$, and $\sigma_{k}=\sigma\leq\frac{2}{L^{\Phi}_{yy}+2\mu}$ and regularization parameter $\mu>0$. Then for any $K\geq 1$ we have that
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\gamma B_{3}D_{\mathcal{X}}+\frac{3}{(1-\rho_{D})(1-\beta)}\frac{L_{y\theta}^{\Phi}}{K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$ (53)
for some constants $B_{1},B_{2},B_{3}>0$ where $B_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $B_{2}=\mathcal{O}\big{(}(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\kappa_{g}^{4}\big{)}$, and $B_{3}\triangleq\mathcal{O}(\frac{\kappa_{g}}{\mu^{2}}+\frac{\kappa_{g}^{2}}{\mu})$.
@@
From Lipschitz continuity of $\nabla f_{\mu}$ we have that
@@
where the last equality follows from adding and subtracting the terms $\langle G_{k}^{x}+\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle$ to RHS. Recall that $x_{k+1}=\gamma s_{k}+(1-\gamma)x_{k}$. Moreover, we define $s^{\prime}_{k}\triangleq\operatorname*{\arg\!\max}_{s\in X}\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s\rangle$ then we have that
@@
$=\gamma\langle\nabla_{x}\mathcal{L}(x_{k},y_{k})_{k},s^{\prime}_{k}-x_{k}\rangle+\gamma\langle G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k}),s^{\prime}_{k}-x_{k}\rangle.$ (55)
Next, using (5.1) within (5.1), rearranging the terms, and using Cauchy Schwartz inequality imply that
@@
$\quad+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}.$ (56)
From Lemma 4.4 we know that $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$. Therefore, using Lipschitz continuity of $\mathcal{L}(x,\cdot)$, $\left\|s^{\prime}_{k}-x_{k}\right\|\leq D_{\mathcal{X}}$, and $\|{x_{k+1}-x_{k}}\|=\gamma\left\|s_{k}-x_{k}\right\|$ we obtain
@@
$\quad+\gamma D_{\mathcal{X}}\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma^{2}\left\|s_{k}-x_{k}\right\|^{2}.$ (57)
Summing the above inequality over $k=0$ to $K-1$ dividing both sides by $\gamma K$, and recalling $\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})=\sup_{s\in\mathcal{X}}\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s\rangle$ imply that for any $K\geq 1$,
@@
From the definition of $f_{\mu}$ and boundedness of $\mathcal{Y}$ we have $f_{\mu}(x_{0})-f_{\mu}(x_{K})\leq f(x_{0})-f(x^{*})+\frac{\mu}{2}D_{\mathcal{Y}}^{2}$. Therefore, for any $K>2$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})$ $\leq\frac{(f(x_{0})-f(x^{*}))}{\gamma K}+\frac{\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+D_{\mathcal{X}}(\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1})}{K}$
$\quad+(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5})D_{\mathcal{X}}^{2}+\gamma\left(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}\right)D_{\mathcal{X}}^{2}+\frac{\mu}{2\gamma K}D_{\mathcal{Y}}^{2}.$
Finally, recall that $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{2}=\mathcal{O}(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma)$, $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}((\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}\gamma+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}\gamma+\kappa_{g}^{4}\gamma)$, $\mathbf{L}_{\mathcal{L}_{1}}=\mathcal{O}(\kappa_{g}^{3})$, and $\mathbf{L}_{\mathcal{L}_{2}}=\mathcal{O}(\kappa_{g})$. Putting the pieces together the desired result can be obtained.
@@
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}\big{(}\left\|y_{k}-\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}\right\|+\left\|\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}-\mathcal{P}_{\mathcal{Y}}(y_{k}+\sigma\nabla_{y}\mathcal{L}(x_{k},y_{k}))\right\|\big{)}.$ (58)
Considering non-expansivity of the projection mapping we have
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}(\left\|y_{k}-y_{k+1}\right\|+\sigma\left\|G_{k}^{y}-\nabla_{y}\mathcal{L}(x_{k},y_{k})\right\|).$ (59)
Based on the conclusion of Lemma 4.9 we can get
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}\left\|y_{k}-y_{k+1}\right\|+L_{y\theta}^{\Phi}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\mu D_{\mathcal{Y}}.$ (60)
Moreover, using Lemma 4.7 and rearragning the terms we conclude that for any $k\geq 0$,
@@
$\quad+\frac{L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\beta\gamma D_{\mathcal{X}}}{1-\beta}+\frac{2L_{y\theta}^{\Phi}}{(1-\rho_{D})(1-\beta)K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}}.$ (62)
Therefore, the desired result can be concluded by noting that $1-\rho_{D}=\mathcal{O}(\mu)$, $1-\beta=\mathcal{O}(1/\kappa_{g})$, and $\mathbf{L}_{\theta}=\mathcal{O}(\kappa_{g})$.
@@
Suppose the premises of Theorem 5.1 holds and we further assume that $L_{yy}^{\Phi}=0$. Let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{3})$ then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$ iterations.

### v3

@@ 4.3 Analysis of Gradients Estimation
In this section, we aim to assess how accurately the gradient estimates, denoted as $G^{x}_{k}$ and $G^{y}_{k}$, approximate the implicit gradients $\nabla_{x}\mathcal{L}(x_{k},y_{k})$ and $\nabla_{y}\mathcal{L}(x_{k},y_{k})$, respectively. Through Lemmas 4.8 and 4.9, we analyze the bounds on the error in estimating these gradients. We demonstrate that the errors may depend linearly on certain terms or involve consecutive terms, helping us in controlling the error. Later, in the convergence analysis, we will also specify the step-sizes to ensure convergence.
@@
where $\Gamma_{1},\Gamma_{3}$ are defined in Lemma 4.6 and 4.7, respectively, $\Gamma_{4}\triangleq D_{\mathcal{X}}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|(\frac{\tilde{C}_{2}}{1-\beta}+\frac{\rho\tilde{C}}{(1-\rho)^{2}})+\frac{\rho C_{\theta x}^{g}D_{\mathcal{X}}}{1-\rho}\left\|w_{0}-v(x_{0},y_{0})\right\|$ and $\Gamma_{5}\triangleq\frac{\gamma\beta\mathbf{L}_{\theta}}{1-\beta}(1+\frac{\tilde{C}}{1-\rho})+\frac{\rho\gamma\mathbf{C}_{v1}C_{\theta x}^{g}}{1-\rho}$. In particular, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$ and $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}\left(\gamma(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\gamma\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\gamma\kappa_{g}^{4}\right)$. Moreover, in case of $L^{\Phi}_{yy}=0$, we have $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}\left(\gamma\frac{\kappa_{g}^{4}}{\mu}(L_{y\theta}^{\Phi})^{2}+\gamma\frac{\kappa_{g}^{2}}{\mu}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\gamma\kappa_{g}^{4}\right)$.
@@
$\quad-\left.\nabla_{\theta x}^{2}g(x_{k},\theta^{*}(x_{k}))v(x_{k},y_{k})\right\|.$
@@
$\leq L_{x\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+C_{\theta x}^{g}\left\|w_{k+1}-v(x_{k},y_{k})\right\|+\frac{C_{\theta}^{\Phi}}{\mu_{g}}L_{\theta x}^{g}\left\|\ \theta_{k}-\theta^{*}(x_{k})\right\|,$
where the last inequality is obtained using Assumption 2.2 and the triangle inequality. Next, utilizing Lemma 4.1 and 4.2, we can further provide upper-bounds for the terms in RHS of the above inequality to obtain the desired result in (4.8).
@@
$\left\|\nabla_{y}\mathcal{L}(x_{k},y_{k})-G_{k}^{y}\right\|\leq\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\mu D_{\mathcal{Y}}.$
@@
$\leq L_{y\theta}^{\Phi}\left\|\theta_{k}-\theta^{*}(x_{k})\right\|+\mu D_{\mathcal{Y}},$
where in the last inequality, we use Lipschitz continuity of $\nabla_{y}\Phi(x,\cdot,y)$ and boundedness of set $\mathcal{Y}$. Next, utilizing Lemma 4.1, we can further provide upper-bounds for the term in RHS of the above inequality as follows
@@ 5.1 Convergence Rate of Algorithm 1
We provide upper bounds for the primal and dual gap functions for the iterates generated by Algorithm 1 in the following theorem. Moreover, we investigate the iteration complexity of i-BRPD:OPF method when the upper-level objective function is linear in $y$ (i.e., $L^{\Phi}_{yy}=0$).
@@
Suppose Assumptions 2.1, 2.2 hold and let $\{(x_{k},y_{k},\theta_{k})\}_{k\geq 0}$ be the sequence generated by Algorithm 1 with step-sizes $\gamma_{k}=\gamma>0$, $\eta_{k}=\eta>0$, $\alpha>0$, and $\sigma_{k}=\sigma\leq\frac{2}{L^{\Phi}_{yy}+2\mu}$ and regularization parameter $\mu>0$. Then, for any $K\geq 1$ we have that
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\gamma B_{3}D_{\mathcal{X}}+\Big{(}1+\frac{2}{1-\rho_{D}}\Big{)}\frac{L_{y\theta}^{\Phi}}{(1-\beta)K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$ (21)
for some constants $B_{1},B_{2},B_{3}>0$ where $B_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $B_{2}=\mathcal{O}\big{(}(\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}+\kappa_{g}^{4}\big{)}$, and $B_{3}=\mathcal{O}(\frac{\kappa_{g}}{\mu^{2}}+\frac{\kappa_{g}^{2}}{\mu})$. Furthermore, if $L^{\Phi}_{yy}=0$, in (21) we have that
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{3}{(1-\beta)}\frac{L_{y\theta}^{\Phi}}{K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma B_{3}D_{\mathcal{X}}+\mu D_{\mathcal{Y}}.$
@@
From Lipschitz continuity of $\nabla f_{\mu}$, we have that
@@
where the last equality follows from adding and subtracting the terms $\langle G_{k}^{x}+\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k+1}-x_{k}\rangle$ to RHS. Recall that $x_{k+1}=\gamma s_{k}+(1-\gamma)x_{k}$. Moreover, we define $s^{\prime}_{k}\triangleq\operatorname*{\arg\!\max}_{s\in X}\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s\rangle$ then, we have that
@@
$=\gamma\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),s^{\prime}_{k}-x_{k}\rangle+\gamma\langle G_{k}^{x}-\nabla_{x}\mathcal{L}(x_{k},y_{k}),s^{\prime}_{k}-x_{k}\rangle.$ (23)
Next, using (5.1) within (5.1), rearranging the terms, and using Cauchy Schwartz inequality implies that
@@
$\quad+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\|{x_{k+1}-x_{k}}\|^{2}.$
From Lemma 4.4, we know that $\nabla f_{\mu}(x)=\nabla_{x}\mathcal{L}(x,y^{*}_{\mu}(x))$. Therefore, using Lipschitz continuity of $\mathcal{L}(x,\cdot)$, $\left\|s^{\prime}_{k}-x_{k}\right\|\leq D_{\mathcal{X}}$, and $\|{x_{k+1}-x_{k}}\|=\gamma\left\|s_{k}-x_{k}\right\|$, we obtain
@@
$\quad+\gamma D_{\mathcal{X}}\|\nabla_{x}\mathcal{L}(x_{k},y_{k})-G_{k}^{x}\|+\frac{1}{2}(\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu})\gamma^{2}\left\|s_{k}-x_{k}\right\|^{2}.$
Summing the above inequality over $k=0$ to $K-1$, dividing both sides by $\gamma K$, and recalling $\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})=\sup_{s\in\mathcal{X}}\langle\nabla_{x}\mathcal{L}(x_{k},y_{k}),x_{k}-s\rangle$ imply that for any $K\geq 1$,
@@
From the definition of $f_{\mu}$ and boundedness of $\mathcal{Y}$, we have $f_{\mu}(x_{0})-f_{\mu}(x_{K})\leq f(x_{0})-f(x^{*})+\frac{\mu}{2}D_{\mathcal{Y}}^{2}$. Therefore, for any $K>2$
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{X}}(x_{k},y_{k})$ $\leq\frac{(f(x_{0})-f(x^{*}))}{\gamma K}+\frac{1}{K}\overbrace{\Big{(}\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{1}+D_{\mathcal{X}}(\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1})\Big{)}}^{\triangleq B_{1}}$
$\quad+\underbrace{\left(\frac{1}{\gamma}(\mathbf{L}_{\mathcal{L}_{2}}\Gamma_{2}+\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5})+\mathbf{L}_{\mathcal{L}_{1}}+\frac{\mathbf{L}_{\mathcal{L}_{2}}L_{y\mu}}{\mu}\right)}_{\triangleq B_{2}}\gamma D_{\mathcal{X}}^{2}+\frac{\mu}{2\gamma K}D_{\mathcal{Y}}^{2}.$
Finally, recall that $\Gamma_{1}=\mathcal{O}(\frac{\kappa_{g}}{\mu}L_{y\theta}^{\Phi})$, $\Gamma_{2}=\mathcal{O}(\frac{\kappa_{g}L_{y\theta}^{\Phi}+L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}}{\mu}\gamma)$, $\Gamma_{3}=\mathcal{O}(\frac{\kappa_{g}^{2}L_{y\theta}^{\Phi}+\kappa_{g}L_{yx}^{\Phi}}{\mu^{2}}\gamma+\frac{\kappa_{g}^{3}L_{y\theta}^{\Phi}}{\mu}\gamma)$, $\Gamma_{4}+2\mathbf{C}_{v2}\Gamma_{1}=\mathcal{O}(\kappa_{g}^{3}+L_{y\theta}^{\Phi}\kappa_{g}^{2}/\mu)$, $\mathbf{C}_{v2}\Gamma_{3}+\Gamma_{5}=\mathcal{O}((\frac{\kappa_{g}^{3}}{\mu^{2}}+\frac{\kappa_{g}^{4}}{\mu})(L_{y\theta}^{\Phi})^{2}\gamma+\frac{\kappa_{g}^{2}}{\mu^{2}}L_{y\theta}^{\Phi}L_{yx}^{\Phi}\gamma+\kappa_{g}^{4}\gamma)$, $\mathbf{L}_{\mathcal{L}_{1}}=\mathcal{O}(\kappa_{g}^{3})$, and $\mathbf{L}_{\mathcal{L}_{2}}=\mathcal{O}(\kappa_{g})$. Putting the pieces together, the desired result in (20) can be obtained.
@@
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}\big{(}\left\|y_{k}-\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}\right\|+\left\|\mathcal{P}_{\mathcal{Y}}\big{(}y_{k}+\sigma G_{k}^{y}\big{)}-\mathcal{P}_{\mathcal{Y}}(y_{k}+\sigma\nabla_{y}\mathcal{L}(x_{k},y_{k}))\right\|\big{)}.$
Considering non-expansivity of the projection mapping, we have
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}(\left\|y_{k}-y_{k+1}\right\|+\sigma\left\|G_{k}^{y}-\nabla_{y}\mathcal{L}(x_{k},y_{k})\right\|).$
Based on the conclusion of Lemma 4.9, we can get
$\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})\leq\frac{1}{\sigma}\left\|y_{k}-y_{k+1}\right\|+L_{y\theta}^{\Phi}\beta^{k}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\gamma L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\sum_{i=0}^{k-1}\beta^{k-i}\left\|s_{i}-x_{i}\right\|+\mu D_{\mathcal{Y}}.$
Moreover, using Lemma 4.7 and rearragning the terms, we conclude that for any $k\geq 0$,
@@
$\quad+\frac{L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\beta\gamma D_{\mathcal{X}}}{1-\beta}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Big{(}1+\frac{2}{1-\rho_{D}}\Big{)}}\frac{L_{y\theta}^{\Phi}}{(1-\beta)K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}},$
$=\underbrace{\Big{(}\frac{2L_{y\mu}\rho_{D}}{\mu(1-\rho_{D})\sigma}+\frac{2L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\beta}{(1-\beta)(1-\rho_{D})}+\frac{L_{y\mu}}{\mu\sigma}+\frac{L_{y\theta}^{\Phi}\mathbf{L}_{\theta}\beta}{1-\beta}\Big{)}}_{\triangleq B_{3}}\gamma D_{\mathcal{X}}$
$\quad+\frac{2\left\|y_{0}-y^{*}_{\mu}(x_{0})\right\|}{\sigma(1-\rho_{D})K}+\Big{(}1+\frac{2}{1-\rho_{D}}\Big{)}\frac{L_{y\theta}^{\Phi}}{(1-\beta)K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\mu D_{\mathcal{Y}}.$ (25)
Therefore, the desired result can be concluded by noting that $1-\rho_{D}=\mathcal{O}(\mu)$, $1-\beta=\mathcal{O}(1/\kappa_{g})$, and $\mathbf{L}_{\theta}=\mathcal{O}(\kappa_{g})$. Furthermore, if $L^{\Phi}_{yy}=0$, then $\rho_{D}=0$, and the bound obtained in (5.1) simplifies to
$\frac{1}{K}\sum_{k=0}^{K-1}\mathcal{G}_{\mathcal{Y}}(x_{k},y_{k})$ $\leq\frac{3}{(1-\beta)}\frac{L_{y\theta}^{\Phi}}{K}\left\|\theta_{0}-\theta^{*}(x_{0})\right\|+\frac{3L^{\Phi}_{y\theta}\mathbf{L}_{\theta}}{(1-\beta)}\gamma D_{\mathcal{X}}+L_{y\mu}\gamma D_{\mathcal{X}}+\mu D_{\mathcal{Y}}.$
@@
Suppose the premises of Theorem 5.1 holds and we further assume that $L_{yy}^{\Phi}=0$. Let $\mu=\mathcal{O}(\epsilon)$, $\gamma=\mathcal{O}(\epsilon^{3})$, then for any $K\geq 1$, there exists $t\in\{0,\ldots,K-1\}$ such that $(x_{t},y_{t})\in\mathcal{X}\times\mathcal{Y}$ satisfies $\mathcal{G}_{\mathcal{Z}}(x_{t},y_{t})\leq\epsilon$ within $\mathcal{O}(\frac{\kappa_{g}^{4}}{\epsilon^{3}})$ iterations.